OpenStudy (anonymous):

Solve: cos( 3/2 θ)*tan(3/2 θ)=0 for 0≤θ≤2

OpenStudy (anonymous):

tan (3/2 θ) = sin (3/2 θ)/ cos (3/2 θ) cos( 3/2 θ)*tan(3/2 θ)=cos( 3/2 θ)*sin (3/2 θ)/ cos (3/2 θ) cos( 3/2 θ)*tan(3/2 θ)=sin (3/2 θ) sin (3/2 θ) = 0 3/2 θ = arcsin 0 + k*pi θ = 2k*pi/3